How do you find the perimeter of a parallelogram when you have the base of 4m, height of 3m and don't have side length?
Answer by drk(1908)

I assume it is isosceles. We can use Pythagorean theorem
a^2 + b^2 = c^2
and then
3^3 + 4^2 = c^2
9 + 16 = c^2
25 = c^2
5 = c.
The perimeter is 3 + 4 + 5 + 5 = 17 m.